# Khoroshkin–Tolstoy approach for quantum superalgebras

### v2

@@ 8. Explicit form of $R$-operator
First construct the operator $K$. Putting $q=1$ in equations (B.4) and (B.5), we obtain expressions for the matrix entries of the matrix $C=B^{-1}$. After some algebra, very similar to that used in the paper [52] for the case of the quantum loop algebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{l+1}))$, we come to the expression
$K=q^{-(M-N-1)/(M-N)}\Big{(}\sum_{i=1}^{M}\mathbb{E}_{ii}\otimes\mathbb{E}_{ii}+q^{2}\sum_{i=M+1}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{ii}+q\sum_{\begin{subarray}{c}i,j,=1\\ i\neq j\end{subarray}}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{jj}\Big{)}.$ (8.1)
@@
$\varphi_{\zeta}(e^{\prime}_{\delta;\,\alpha_{i}})=-\zeta^{s}(-1)^{M-1}q^{M}(\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
$\varphi_{\zeta}(f^{\prime}_{\delta;\,\alpha_{i}})=\zeta^{-s}(-1)^{-M+1}q^{-M}(\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
@@
$\varphi_{\zeta}(e_{(\delta-\alpha_{ij})+n\delta})=\zeta^{(s-s_{i,\,i+1})+ns}(-1)^{(n+1)i+n}q^{(n+1)i+n}\mathbb{E}_{i+1,\,i},$ (8.10)
@@
$\varphi_{\zeta}(e_{(\delta-\alpha_{ij})+n\delta})=\zeta^{(s-s_{ij})+ns}(-1)^{(n+1)i+n}q^{(n+1)i}\mathbb{E}_{ji},$ (8.12)
@@
$\varphi_{\zeta}(e_{(\delta-\alpha_{ij})+n\delta})=-\zeta^{(s-s_{ij})+ns}(-1)^{(n+1)M}q^{(n+1)M+n}\mathbb{E}_{ji}$ (8.14)
$\varphi_{\zeta}(f_{(\delta-\alpha_{ij})+n\delta})=\zeta^{-(s-s_{ij})-ns}(-1)^{-(n+1)M-n+1}q^{-(n+1)M-n}\mathbb{E}_{ij},$ (8.15)
@@
$\varphi_{\zeta}(e_{(\delta\alpha_{ij})+n\delta})=-\zeta^{(s-s_{ij})+ns}(-1)^{(n+1)i+1}q^{(n+1)(2M-i+1)+1}\mathbb{E}_{ji},$ (8.16)
$\varphi_{\zeta}(f_{(\delta-\alpha_{ij})+n\delta})=\zeta^{-(s-s_{ij})-ns}(-1)^{-(n+1)i-n}q^{-(n+1)(2M-i+1)-1}\mathbb{E}_{ij},$ (8.17)
@@
$\varphi_{\zeta}(e^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{ns}(-1)^{ni+n+1}q^{n(i+1)-1}\big{(}\mathbb{E}_{ii}-q^{2}\mathbb{E}_{i+1,\,i+1}\big{)},$ (8.19)
$\varphi_{\zeta}(f^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni-1}q^{-n(i+1)+1}\big{(}\mathbb{E}_{ii}-q^{-2}\mathbb{E}_{i+1,\,i+1}\big{)},$ (8.20)
@@
$\varphi_{\zeta}(e^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{ns}(-1)^{nM+1}q^{n(M+1)-1}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$ (8.21)
$\varphi_{\zeta}(f^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-nM-n-1}q^{-n(M+1)+1}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$ (8.22)
@@
$\varphi_{\zeta}(e^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{ns}(-1)^{ni+1}q^{n(2M-i+1)+1}\big{(}\mathbb{E}_{ii}-q^{-2}\mathbb{E}_{i+1,\,i+1}\big{)},$ (8.23)
$\varphi_{\zeta}(f^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni-n-1}q^{-n(2M-i+1)-1}\big{(}\mathbb{E}_{ii}-q^{2}\mathbb{E}_{i+1,\,i+1}\big{)}.$ (8.24)
Now we find expressions for $\varphi_{\zeta}(e_{n\delta,\,\alpha_{ij}})$ and $\varphi_{\zeta}(f_{n\delta,\,\alpha_{ij}})$. Using (8.19)–(8.24), we obtain for the corresponding generating functions the following expressions
@@
$\mathbbm{1}=\sum_{i,j=1}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{jj},$
and come to the equations
@@
$\varphi_{\zeta}(e_{n\delta;\,\alpha_{i}})=\zeta^{ns}(-1)^{ni+n+1}q^{ni}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{2n}\mathbb{E}_{i+1,\,i+1}\big{)},$
$\varphi_{\zeta}(f_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni-1}q^{-ni}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{-2n}\mathbb{E}_{i+1,\,i+1}\big{)},$
@@
$\varphi_{\zeta}(e_{n\delta;\,\alpha_{i}})=\zeta^{ns}(-1)^{nM+1}q^{nM}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
$\varphi_{\zeta}(f_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-nM-n-1}q^{-nM}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
@@
$\varphi_{\zeta}(e_{n\delta;\,\alpha_{i}})=\zeta^{ns}(-1)^{ni+1}q^{n(2M-i+2)}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{-2n}\mathbb{E}_{i+1,\,i+1}\big{)},$
$\varphi_{\zeta}(f_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni-n-1}q^{-n(2M-i+2)}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{2n}\mathbb{E}_{i+1,\,i+1}\big{)}$
@@
$T_{nij}=\frac{[nB_{ij}]_{q}}{n}=\frac{[n]_{q}}{n}[B_{ij}]_{q^{n}},$
where $B_{ij}$ are the matrix entries of the symmetrized Cartan matrix $B$ of the Lie superalgebra $\mathfrak{sl}_{M|N}$. It is clear that to find the inverse of $T_{n}$ it is sufficient to find the inverse of the matrix
$B_{q}=([B_{ij}]_{q})_{i,j=1}^{M+N-1}.$
@@
and performing some summations, we obtain
@@
$\hskip 40.00006pt{}=-(F_{M-N}(q^{M-N-1}\zeta^{s}_{12})-F_{M-N}(q^{-(M-N-1)}\zeta^{s}_{12}))\sum_{i,\,j=1}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{jj}$
@@
Using equations (8.3) and (8.8), we determine that
$(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\alpha_{ij}+n\delta})=\exp_{(-1)^{[i]+[j]}q^{d_{i}+d_{j}}}\big{(}-\zeta_{12}^{s_{ij}+ns}(q-q^{-1})(-1)^{[j]}(\mathbb{E}_{ij}\otimes\mathbb{E}_{ji})\big{)}.$
@@
for all $1\leq i<j\leq M+N$ and $1\leq k<m\leq M+N$, we see that the factors $(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\alpha_{ij}+n\delta})$ of $(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\prec\delta})$ can be taken in an arbitrary order and obtain the expression
@@ 9. Conclusion
In this paper, we have studied a generalization of the Khoroshkin–Tolstoy approach to the construction of a universal $R$-matrix, which is necessary in the case of quantum superalgebras. In fact, our consideration can be seen as a refinement and completion of the general procedure outlined in the papers [24, 25, 26]. Using the obtained results, we have constructed an explicit expression for the $R$-operator for the case of a vector evaluation representation of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. In the future, we plan to construct expressions for other integrability objects and obtain the corresponding functional relations for them.
@@ Acknowledgments
This work was supported in part by the RFBR grant # 20-51-12005. The author is grateful to H.  Boos, F.  Göhmann, A.  Klümper, and Kh. S.  Nirov, in collaboration with whom some important results, used in this paper, were previously obtained, for useful discussions.
@@ Appendix A Superalgebras and Lie superalgebras
a superspace. The elements of the subspaces $V_{\overline{0}}$ and $V_{\overline{1}}$ are said to be even and odd, respectively. An element belonging to $V_{\overline{0}}$ or $V_{\overline{1}}$ is said to be homogeneous. The parity of a homogeneous element $v$, denoted as $[v]$, is $\overline{0}$ or $\overline{1}$ according to whether it is in $V_{\overline{0}}$ or $V_{\overline{1}}$. [^7] Throughout the paper we use the following convenient convention. If $v$ is an element of a super vector space and $[v]$ appears in some formula or expression, then $v$ is assumed to be homogeneous. For a finite dimensional superspace $V$, the pair $(\dim V_{\overline{0}},\,\dim V_{\overline{1}})$ is called the dimension of $V$. The tensor product $V\otimes W$ of two superspaces $V$ and $W$ is considered as a superspace with
@@
$A_{\alpha}A_{\beta}\subset A_{\alpha+\beta},\qquad\alpha,\beta={\overline{0}},{\overline{1}}.$
@@
A $\mathbb{Z}_{2}$-graded associative algebra is said to be an associative superalgebra. The tensor product $A\otimes B$ of two associative superalgebra $A$ and $B$ is considered as an associative superalgebra with the multiplication defined by the equation
$(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})=(-1)^{[b_{1}][a_{2}]}(a_{1}a_{2})\otimes(b_{1}b_{2})$
for all homogeneous $a_{1},a_{2}\in A$ and $b_{1},b_{2}\in B$. The multiplication in the tensor product of any number of associative superalgebras id defined recursively.
Let $V$ be a superspace. The algebra $\mathrm{End}(V))$ becomes an associative superalgebra if one assumes that $a\in(\mathrm{End}(V))_{\overline{0}}$ when $aV_{\alpha}\subset V_{\alpha}$, and $a\in(\mathrm{End}(V)_{\overline{1}}$ when $aV_{\alpha}\subset V_{\alpha+{\overline{1}}}$ for $\alpha={\overline{0}},{\overline{1}}$.
@@
$(-1)^{[x][z]}[x,\,[y,\,z]]+(-1)^{[y][x]}[y,\,[z,\,x]]+(-1)^{[z][y]}[z,\,[x,\,y]]=0$
for any homogeneous elements $z,y,z\in\mathfrak{g}$. Here and throughout the article, it is assumed that

### v3

@@ 8. Explicit form of $R$-operator
First, construct the expression for the operator $K_{\varphi}$. Putting $q=1$ in equations (B.4) and (B.5), we obtain expressions for the matrix entries of the matrix $C=B^{-1}$. Now, using equations (7.6), (6.7), and (6.6), we get
$K_{\varphi}\,v_{m}\otimes v_{n}=q^{-\sum_{i,\,j=1}^{M+N-1}d_{i}^{-1}d_{j}^{-1}\langle\omega_{1}-\sum_{k=1}^{m-1}\alpha_{k},\,h_{i}\rangle\langle\omega_{1}-\sum_{l=1}^{n-1}\alpha_{l},\,h_{j}\rangle c_{ij}}\,v_{m}\otimes v_{n}\\ =q^{-(d_{1}^{-2}c_{11}-d_{1}^{-1}\sum_{k=1}^{m-1}\delta_{1k}-d_{1}^{-1}\sum_{l=1}^{n-1}\delta_{l1}+\sum_{k=1}^{m-1}\sum_{l=1}^{n-1}b_{kl})}\,v_{m}\otimes v_{n}.$
One can demonstrate that
$K_{\varphi}\,v_{m}\otimes v_{m}=q^{-(M-N-1)/(M-N)}v_{m}\otimes v_{m},$ $m\leq M,$
$K_{\varphi}\,v_{m}\otimes v_{m}=q^{-(M-N-1)/(M-N)+2}\,v_{m}\otimes v_{m},$ $m>M,$
$K_{\varphi}\,v_{m}\otimes v_{n}=q^{-(M-N-1)/(M-N)+1}\,v_{m}\otimes v_{n}.$ $m\neq n.$
Using these equations, we obtain the expression
$K_{\varphi}=q^{-(M-N-1)/(M-N)}\Big{(}\sum_{i=1}^{M}\mathbb{E}_{ii}\otimes\mathbb{E}_{ii}+q^{2}\sum_{i=M+1}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{ii}+q\sum_{\begin{subarray}{c}i,j,=1\\ i\neq j\end{subarray}}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{jj}\Big{)}.$ (8.1)
@@
$\varphi_{\zeta}(e^{\prime}_{\delta;\,\alpha_{i}})=-\zeta^{s}(-1)^{M}q^{M}(\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
$\varphi_{\zeta}(f^{\prime}_{\delta;\,\alpha_{i}})=\zeta^{-s}(-1)^{-M}q^{-M}(\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
@@
$\varphi_{\zeta}(e_{(\delta-\alpha_{ij})+n\delta})=-\zeta^{(s-s_{i,\,i+1})+ns}(-1)^{(n+1)i+n+1}q^{(n+1)i+n}\mathbb{E}_{i+1,\,i},$ (8.10)
@@
$\varphi_{\zeta}(e_{(\delta-\alpha_{ij})+n\delta})=-\zeta^{(s-s_{ij})+ns}(-1)^{(n+1)i+n+1}q^{(n+1)i}\mathbb{E}_{ji},$ (8.12)
@@
$\varphi_{\zeta}(e_{(\delta-\alpha_{ij})+n\delta})=-\zeta^{(s-s_{ij})+ns}(-1)^{(n+1)M+1}q^{(n+1)M+n}\mathbb{E}_{ji}$ (8.14)
$\varphi_{\zeta}(f_{(\delta-\alpha_{ij})+n\delta})=\zeta^{-(s-s_{ij})-ns}(-1)^{-(n+1)M+n+1}q^{-(n+1)M-n}\mathbb{E}_{ij},$ (8.15)
@@
$\varphi_{\zeta}(e_{(\delta\alpha_{ij})+n\delta})=-\zeta^{(s-s_{ij})+ns}(-1)^{(n+1)i}q^{(n+1)(2M-i+1)+1}\mathbb{E}_{ji},$ (8.16)
$\varphi_{\zeta}(f_{(\delta-\alpha_{ij})+n\delta})=\zeta^{-(s-s_{ij})-ns}(-1)^{-(n+1)i+n}q^{-(n+1)(2M-i+1)-1}\mathbb{E}_{ij},$ (8.17)
@@
$\varphi_{\zeta}(e^{\prime}_{n\delta;\,\alpha_{i}})=-\zeta^{ns}(-1)^{ni+n}q^{n(i+1)-1}\big{(}\mathbb{E}_{ii}-q^{2}\mathbb{E}_{i+1,\,i+1}\big{)},$ (8.19)
$\varphi_{\zeta}(f^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni+1}q^{-n(i+1)+1}\big{(}\mathbb{E}_{ii}-q^{-2}\mathbb{E}_{i+1,\,i+1}\big{)},$ (8.20)
@@
$\varphi_{\zeta}(e^{\prime}_{n\delta;\,\alpha_{i}})=-\zeta^{ns}(-1)^{nM}q^{n(M+1)-1}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$ (8.21)
$\varphi_{\zeta}(f^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-nM+n+1}q^{-n(M+1)+1}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$ (8.22)
@@
$\varphi_{\zeta}(e^{\prime}_{n\delta;\,\alpha_{i}})=-\zeta^{ns}(-1)^{ni}q^{n(2M-i+1)+1}\big{(}\mathbb{E}_{ii}-q^{-2}\mathbb{E}_{i+1,\,i+1}\big{)},$ (8.23)
$\varphi_{\zeta}(f^{\prime}_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni+n+1}q^{-n(2M-i+1)-1}\big{(}\mathbb{E}_{ii}-q^{2}\mathbb{E}_{i+1,\,i+1}\big{)}.$ (8.24)
Now, we find expressions for $\varphi_{\zeta}(e_{n\delta,\,\alpha_{ij}})$ and $\varphi_{\zeta}(f_{n\delta,\,\alpha_{ij}})$. Using (8.19)–(8.24), we obtain for the corresponding generating functions the following expressions
@@
$\mathbbm{1}=\sum_{i,j=1}^{M+N}\mathbb{E}_{ii}\otimes\mathbb{E}_{jj}$
is the identity element of the algebra $\mathrm{End}(\mathbb{C}_{M|N})$, and come to the equations
@@
$\varphi_{\zeta}(e_{n\delta;\,\alpha_{i}})=-\zeta^{ns}(-1)^{ni+n}q^{ni}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{2n}\mathbb{E}_{i+1,\,i+1}\big{)},$
$\varphi_{\zeta}(f_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni+1}q^{-ni}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{-2n}\mathbb{E}_{i+1,\,i+1}\big{)},$
@@
$\varphi_{\zeta}(e_{n\delta;\,\alpha_{i}})=-\zeta^{ns}(-1)^{nM}q^{nM}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
$\varphi_{\zeta}(f_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-nM+n+1}q^{-nM}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{MM}+\mathbb{E}_{M+1,\,M+1}\big{)},$
@@
$\varphi_{\zeta}(e_{n\delta;\,\alpha_{i}})=-\zeta^{ns}(-1)^{ni}q^{n(2M-i+2)}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{-2n}\mathbb{E}_{i+1,\,i+1}\big{)},$
$\varphi_{\zeta}(f_{n\delta;\,\alpha_{i}})=\zeta^{-ns}(-1)^{-ni+n+1}q^{-n(2M-i+2)}\frac{[n]_{q}}{n}\big{(}\mathbb{E}_{ii}-q^{2n}\mathbb{E}_{i+1,\,i+1}\big{)}$
@@
$t_{nij}=o_{i}^{n}o_{j}^{n}\frac{[nb_{ij}]_{q}}{n}=o_{i}^{n}o_{j}^{n}\frac{[n]_{q}}{n}[b_{ij}]_{q^{n}}.$
It is clear that to find the inverse of $T_{n}$ it is sufficient to find the inverse of the matrix
$B_{q}=([b_{ij}]_{q})_{i,j=1}^{M+N-1}.$ (8.25)
@@
and performing the summations over $n$, we obtain
@@
$\hskip 40.00006pt{}=-(F_{M-N}(q^{M-N-1}\zeta^{s}_{12})-F_{M-N}(q^{-(M-N-1)}\zeta^{s}_{12}))\sum_{i,\,j=1}^{M-N}\mathbb{E}_{ii}\otimes\mathbb{E}_{jj}$
@@
Using equations (8.3) and (8.8), we find that
$(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\alpha_{ij}+n\delta})=\exp_{(-1)^{[i]+[j]}q_{i}q_{j}}\big{(}-\zeta_{12}^{s_{ij}+ns}(q-q^{-1})(-1)^{[j]}(\mathbb{E}_{ij}\otimes\mathbb{E}_{ji})\big{)}.$
@@
for all $1\leq i<j\leq M+N$ and $1\leq k<m\leq M+N$, we see that the factors $(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\alpha_{ij}+n\delta})$ of the operator $(\varphi_{\zeta_{1}}\otimes\varphi_{\zeta_{2}})(\mathcal{R}_{\prec\delta})$ can be taken in an arbitrary order and obtain the expression
@@ 9. Conclusion
In this paper, we have studied a generalization of the Khoroshkin–Tolstoy approach to the construction of a universal $R$-matrix, which is necessary in the case of quantum superalgebras. In fact, our consideration can be seen as a refinement of the general procedure outlined in the papers [24, 25, 26]. Using the obtained results, we have constructed an explicit expression for the $R$-operator for the case of a vector evaluation representation of the quantum superalgebra $\mathrm{U}_{q}(\mathcal{L}(\mathfrak{sl}_{M|N}))$. In the future, we plan to construct expressions for other integrability objects and obtain for them the corresponding functional relations.
@@ Acknowledgments
This work was supported in part by the RFBR grant # 20-51-12005. The author is grateful to H.  Boos, F.  Göhmann, A.  Klümper, and Kh. S. Nirov in collaboration with whom some important results, used in this paper, were previously obtained, for useful discussions.
Conflict of interest
The author declare that he has no conflicts of interest.
@@ Appendix A Superalgebras and Lie superalgebras
a superspace. The elements of the subspaces $V_{\overline{0}}$ and $V_{\overline{1}}$ are said to be even and odd, respectively. An element belonging to $V_{\overline{0}}$ or $V_{\overline{1}}$ is said to be homogeneous. The parity of a homogeneous element $v$, denoted as $[v]$, is $\overline{0}$ or $\overline{1}$ according to whether it is in $V_{\overline{0}}$ or $V_{\overline{1}}$. [^8] Throughout the paper we use the following convenient convention. If $v$ is an element of a superspace and $[v]$ appears in some formula or expression, then $v$ is assumed to be homogeneous. For a finite dimensional superspace $V$, the pair $(\dim V_{\overline{0}},\,\dim V_{\overline{1}})$ is called the dimension of $V$. The tensor product $V\otimes W$ of two superspaces $V$ and $W$ is considered as a superspace with
@@
$A_{\alpha}A_{\beta}\subset A_{\alpha+\beta},\qquad\alpha,\beta\in\mathbb{Z}_{2}.$
@@
A superalgebra with an associative multiplication is called an associative superalgebra. The tensor product $A\otimes B$ of two associative superalgebras $A$ and $B$ is considered as an associative superalgebra with the multiplication defined by the equation
$(a_{1}\otimes b_{1})(a_{2}\otimes b_{2})=(-1)^{[b_{1}][a_{2}]}(a_{1}a_{2})\otimes(b_{1}b_{2}).$
The multiplication in the tensor product of any number of associative superalgebras is defined recursively.
Let $V$ be a superspace. The algebra $\mathrm{End}(V))$ becomes an associative superalgebra if one assumes that $a\in(\mathrm{End}(V))_{\overline{0}}$ when $aV_{\alpha}\subset V_{\alpha}$, and $a\in(\mathrm{End}(V)_{\overline{1}}$ when $aV_{\alpha}\subset V_{\alpha+{\overline{1}}}$ for $\alpha\in\mathbb{Z}_{2}$.
@@
$(-1)^{[x][z]}[x,\,[y,\,z]]+(-1)^{[y][x]}[y,\,[z,\,x]]+(-1)^{[z][y]}[z,\,[x,\,y]]=0.$
Here and throughout the article, it is assumed that